there are 29 students, the number of girls in class is 7 less than two times the number of boys. how many of the students are girls and how many are boys?
thank you so much! :)
sincerely, emily

Answer by scott8148(6628)

"there are 29 students" ___ g + b = 29

"the number of girls in class is 7 less than two times the number of boys" ___ g = 2b - 7

substituting ___ (2b - 7) + b = 29 ___ 3b - 7 = 29 ___ 3b = 36 ___ b = 12

substituting ___ g + (12) = 29 ___ g = 17